OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

Since it has the same slope, get rid of the y-intercept and solve agin. y = x + 7 -2 = 6 + b b = -8 y = x - 8

OpenStudy (anonymous):

or you can use point slope form. y-y1=m(x-x1) where (x1,y1) is a point in the line and m=slope We know the slope is 1 right since it's the same as y=x+7 Ok now we have : y-y1=x-x1 y+2=x-6 Solve for y